OpenStudy (anonymous):

^_^ Alaine has 1 gallon of paint. She is going to pour it into a paint tray that measures 10 inches wide, 12 inches long, and 5 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? a. The paint will not fill the tray by 369 cm3. b. The paint will not fill the tray by 5.22 in3. c. The paint will fill the tray exactly. d. The paint will overfill the tray by 5.22 in3

OpenStudy (anonymous):

first do you know how to find the volume of a rectangle? you need to find that first because in^3 is the volume. if it was in^2 then it would just be area.

OpenStudy (anonymous):

oh can you help me pleaseee n youll be my best friend forevr im not so good at math

OpenStudy (anonymous):

the volume for a rectangle is V=l*w*h so V=(10*12*5) and what do you get?

OpenStudy (anonymous):

oh ok so i just do l x w x h and there is my answer?

OpenStudy (anonymous):

no. your answer will be B

OpenStudy (anonymous):

it will not fill the tray by 5.22in3

OpenStudy (anonymous):

ohh well i kinda get it but thanks so let me get this straight? i did lxwxh and then i make it in to the inches^3 then from there see how much cm it is?

OpenStudy (anonymous):

231×2.543=3785.41 10×12×1.968=236.22in 325.4×30.48×5=3870.96cm3 she's about to pour 231in3 or 3785.41cm3 well, obviously the tray can take more than 231 cubic inches, so it will fit notice that, the tray can hold 236.22 the paint is just 231 the difference is 5.22 the tray holds the paint, and still have more room for 5.22 cubic inches so The paint will not fill the tray by 5.22 in3.

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yesh thanksss ^_^

OpenStudy (anonymous):

no problem(: